HETAV PATEL

Grade 11,

30sin+90cos=

30sin+90cos=

Grade:10

13 Answers

Vijay Mukati
8 years ago
Please also write the value of angle theta.
runo
8 years ago
30sin+90cos=1/2 + 0=1/2
therefore,30sin +90cos=1/2
Prajwal Kavad
8 years ago
sin 30 + cos 90= 1/2
KARTHIK
8 years ago
0.5
 
venky
8 years ago
1/2
OUTPASS
8 years ago
0.5
RAJU
8 years ago
Ask question properly
saibhargav
8 years ago
½
saibhargav
8 years ago
ASK QUESTION PROPERLY
 
Raghu Vamshi Hemadri
8 years ago
ASK QUESTION PROPERLY
SHANMUKESHWAR
8 years ago
in compleat question
vamshi
8 years ago
1/2
SEELA YUVARAJU
8 years ago
1/2